# A Universal Scheme for Dynamic Partitioned Shortest Path Index

### v2

@@ Abstract.
Shortest path (SP) computation is the fundamental operation in various networks such as urban networks, logistic networks, communication networks, social networks, etc. With the development of technology and societal expansions, those networks tend to be massive. This, in turn, causes deteriorated performance of SP computation, and graph partitioning is commonly leveraged to scale up the SP algorithms. However, the partitioned shortest path (PSP) index has never been systematically investigated and theoretically analyzed, and there is a lack of experimental comparison among different PSP indexes. Moreover, few studies have explored PSP index maintenance in dynamic networks. Therefore, in this paper, we systematically analyze the dynamic PSP index by proposing a universal scheme for it. Specifically, we first propose two novel partitioned shortest path strategies (No-boundary and Post-boundary strategies) to improve the performance of PSP indexes and design the corresponding index maintenance approaches to deal with dynamic scenarios. Then we categorize the partition methods from the perspective of partition structure to facilitate the selection of partition methods in the PSP index. Furthermore, we propose a universal scheme for designing the PSP index by coupling its three dimensions (i.e. PSP strategy, partition structure, and SP algorithm). Based on this scheme, we propose five new PSP indexes with prominent performance in either query or update efficiency. Lastly, extensive experiments are implemented to demonstrate the effectiveness of the proposed PSP scheme, with valuable guidance provided on the PSP index design.
PVLDB Reference Format:
PVLDB, (): , .
doi: This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. , No.  ISSN 2150-8097.
doi:
footnotetext: *Both authors contributed equally to this paper. Lei Li is the corresponding author.
@@ 1. Introduction
Shortest Path (SP) query in dynamic networks is an essential building block for various applications affecting daily lives. Given an origin-destination pair, it returns the minimum traveling time in a road network (Bast et al., 2016), the fastest connection in the web graph (Boccaletti et al., 2006), or the most intimate relationships in the social network (Kolaczyk et al., 2009; Opsahl et al., 2010; Yao et al., 2013). With the development of urban traffic systems and the evolvement of online interactions platforms, real-life networks tend to be massive, which brings great challenges to the scalability of the state-of-the-art (Li et al., 2020a, 2019b; Zhang et al., 2021a, c, b, 2023a) with either heavy memory or expensive search overheads. For example, 2-Hop Labeling (Cohen et al., 2003) has the fastest query efficiency but requires $O(nm^{1/2})$ space (for a graph with $n$ vertices and $m$ edges) to store the index. Complementary, the search-based methods like Dijkstra’s are inefficient on query processing on large networks. Moreover, because the networks are dynamic in nature with evolving structures or edge weights, extra efforts are required to handle the updates for SP indexes.
@@
Graph partitioning is often used to improve the scalability of SP algorithms, enabling more complicated problems like time-dependent (Li et al., 2019c, 2020b) and constraint path computation (Wang et al., 2016; Liu et al., 2021, 2022a) to be efficiently handled in large networks. It decomposes a large network into several smaller ones such that the construction time can be improved by parallelization, and the index sizes and update workloads can be reduced. As shown in Figure 1, we can compare them briefly in terms of index construction time, storage space, query time, and update time. On one extreme are the direct search algorithms (Dijkstra, 1959; Hart et al., 1968; Li et al., 2020c; Zhang et al., 2020) that require no index and can work in dynamic environments directly but are slow for query answering. Then the partitioned search adds various information to guide and reduce search space (Chondrogiannis and Gamper, 2016; Möhring et al., 2007; Bauer and Delling, 2010) to improve query processing. On the other extreme, Contraction Hierarchy (CH) (Geisberger et al., 2008) and 2-hop Labeling (HL) (Cohen et al., 2003; Akiba et al., 2013; Li et al., 2019b; Ouyang et al., 2018) have larger index sizes, longer construction and maintenance time but faster query performance. Their partitioned versions (Delling et al., 2017; Li et al., 2020a; Li et al., 2019c; Wang et al., 2021; Farhan et al., 2022; Liu et al., 2021, 2022b, 2022a) are faster to construct with smaller sizes but longer query times. As these works have better performance than their unpartitioned versions, they create the following misconceptions regarding the benefits of partitions on pathfinding-related problems: 1) Graph partitioning and shortest path algorithms are irrelevant since they are two different research branches; 2) Different PSP indexes are irrelevant and very different from each other; 3) Applying graph partitioning is always better.
@@
Motivations. These misconceptions are not always true. Although the partitioned shortest path (PSP) indexes have been widely applied in the past decade, to the best of our knowledge, they have not been studied systematically. Specifically, the existing solutions typically pick one partition method without discriminating their characteristics and figure out a way to make it work, and then claim benefits/superiority. Consequently, there lacks a generalized scheme to organize and compare the PSP indexes insightfully and fairly. Towards that, we postulate that both theoretical and experimental systematic study is needed for PSP indexes, so that a suitable index could be constructed for different scenarios.
Challenges. However, it is non-trivial to design such a scheme. Firstly, the current PSP indexes are delicately designed structures with tightly coupled components and all claim certain superiority. However, since no effort has been put into the abstract PSP index itself, it is unclear what determines their performance, when to use particular indexes, and how to design a suitable new PSP index for a new scenario. Therefore, we study the PSP index systematically in this work and identify three dimensions throughout the rich literature. Secondly, it is unclear how to construct, query, and update a new PSP index as all existing works only claim their unique structures can answer queries correctly and seldom involve index maintenance. Therefore, we dive into the partitioned shortest path (PSP) strategy, which determines how to coordinate multiple partitions for correct query answering, and summarizes the traditional PSP strategy. Additionally, we propose two novel PSP strategies together with their index update methods to improve the index performance. Besides, a pruning-based overlay graph optimization is also proposed to improve the index and update efficiency further. Thirdly, the graph partition methods are normally classified from different perspectives (like vertex-cut / edge-cut, in-memory / streaming, etc.), but these criteria are not path-oriented and thus hardly help to tell which partition method is suitable for certain path-oriented applications. Therefore, we propose a path-oriented partitioning classification from the perspective of partition structure for easier partition method selection. Lastly, after elaborating on each dimension of PSP index, we propose our universal PSP index scheme by entangling these dimensions. Based on this scheme, we propose five new PSP indexes to achieve better efficiency in either query or update operations by selecting the specific strategy or method on each dimension as per system requirements.
Contributions. Our contributions are listed as follows:
@@
    We propose a universal PSP index scheme by decoupling the PSP index into three dimensions: PSP strategy, partition structure, and SP algorithm, such that the PSP index can be analyzed and designed systematically. (Section 2)
@@
    We propose two novel PSP strategies called No-Boundary and Post-Boundary strategies to improve the index construction and maintenance efficiency over the traditional PSP strategy. We also provide non-trivial correctness analysis for their index construction, query processing, and index maintenance. A novel Pruning-based overlay graph optimization is also designed to prune the unnecessary computation. (Section 3)
@@
    We propose a new path-oriented partitioning classification from the perspective of partition structure for easier partition method selection in designing the PSP index. (Section 4)
@@
    Based on the proposed PSP index scheme, we put forward a PSP index generator to create new PSP structures. Among them, we identify and design five new PSP indexes suitable for specific network structures and scenarios. (Section 5)
  • $\bullet$
@@
    We implement comprehensive experimental studies to explore the performance of PSP indexes under our proposed universal scheme and provide guidelines for designing PSP indexes with different application scenarios. (Section 6)
2. Preliminary and PSP Dimensions
@@
In this paper, we focus on the dynamic weighted network $G(V,E,W)$ with $V$ denoting the vertex set, $E$ denoting the edge set and $W:E\rightarrow\mathbf{R}^{+}$ assigning a non-negative weight to each edge – i.e. $w(u,v)\in W,\forall(u,v)\in E$ – which can increase or decrease in an ad-hoc manner. We denote the number of vertices and edges as $n=|V|$ and $m=|E|$. For each $v\in V$, we represent its neighbors as $N(v)=\{u,|(u,v)\in E\}$, u’s commonly called adjacent to $v$. We associate each vertex $v\in V$ with order $r(v)$ indicating its importance in $G$. In this paper, we use boundary-first ordering (Liu et al., 2022a) to generate the order for the PSP indexes. For non-partitioned SP indexes, we leverage degree to decide the order of small-world networks while the minimum degree elimination-based ordering (Berry et al., 2003) for the road networks. A path $p=\left<v_{0},v_{1},\dots,v_{k}\right>((v_{i},v_{i+1})\in E,0\leq i<k)$ is a sequence of adjacent vertices with length of $l(p)=\sum_{i=0}^{k-1}w(v_{i},v_{i+1})$. Given an origin $s$ and destination $t$ vertex (OD pair), the shortest path query $Q(s,t)$ calculates the path with the minimum length between them, i.e. $p_{G}(s,t)$ with the shortest distance $d_{G}(s,t)$. In the following, we introduce the PSP index dimensions and summarize each of them.
2.1. Dimensions of PSP Index
Based on a comprehensive literature review and analysis of the PSP index, we identify three critical dimensions of the PSP index: 1) Graph Partition Method, which divides the graph into multiple subgraphs and is the key difference from the standalone SP indexes. However, the existing partition methods are not designed for path indexes, so how they affect the PSP index is unknown and deserves study; 2) PSP Strategy, which provides a procedure to organize the index construction for multiple subgraphs and coordinate those indexes such that the query correctness is guaranteed. Only straightforward solutions exist currently as will be discussed in Section 2.3. However, this part is crucial because it determines the complexity of both the query and update; 3) SP Algorithm, whose influence was thoroughly studied in our previous work (Zhang et al., 2021a). With these three dimensions, we can form a space to position the existing methods into it. Given a massive dynamic network, these dimensions provide a roadmap to construct a PSP index by selecting a specific graph partition method, SP algorithm and PSP strategy such that the application requirement (query/update efficiency) can be satisfied. Next, we discuss and review these three dimensions.
2.2. Graph Partitioning
Definition 0 (Graph Partitioning).
A graph $G$ is decomposed into multiple disjoint subgraphs $\{G_{i}\}$ $(i=1,2,\dots,k)$ with $V_{i}\cap V_{j}=\emptyset$ and $\bigcup V_{i}=V$. denoted as $E_{inter}$ and $E_{intra}$.
We now discuss and categorize the vertex and edge in graph partitioning. $\forall v\in G_{i}$, we say $v$ is a boundary vertex if there exists a neighbor of $v$ in the another subgraph, that is $\exists u\in N(v),u\in G_{j}(i\neq j)$. Otherwise, $v$ is an inner vertex. We represent the boundary vertex set of $G_{i}$ as $B_{i}$ and that of $G$ as $B=\bigcup B_{i}$. For $(u,v)\in E$, we say $(u,v)$ is an inter-edge if both its two endpoints $u$ and $v$ are boundary vertices from different subgraphs, i.e., $u\in B_{i},v\in B_{j},i\neq j$. Otherwise, it is an intra-edge. The corresponding edge sets are denoted as $E_{inter}$ and $E_{intra}$, respectively. For example in Figure 2, the example graph is partitioned into four subgraphs $G_{1}$ to $G_{4}$, with the red vertices denoting the boundaries ($B_{1}=\{v_{3},v_{10}\}$) and the white vertices representing the inner vertices. Edges outside partitions like $(v_{3},v_{5})$ are inter-edges, and $(v_{3},v_{10})$ is an intra-edge. Besides, the boundary vertex plays an important role in the shortest path computation with its cut property: given a shortest path $p(s,t)$ with its endpoints in two different subgraphs $s\in G_{i},t\notin G_{i}$, there is at least one vertex $v$ which is in the boundary vertex set of $G_{i}$ on the path, that is $\exists v\in p(s,t),v\in B_{i}$. Therefore, boundary vertices would determine the correctness and complexity of partitioned shortest path computation (discussed in the following sections). Various partition methods are involved in the existing PSP methods such as METIS (Karypis and Kumar, 1998) in G-Tree (Zhong et al., 2013) and SHARC (Bauer and Delling, 2010) ; PUNCH (Delling et al., 2011) in COLA (Wang et al., 2016), FHL (Liu et al., 2021, 2022b), FHL-Cube (Liu et al., 2022a) and THop (Li et al., 2019c, 2020b) ; Tree Decomposition-based Partition (Ouyang et al., 2018) in Core-Tree (Li et al., 2020a, 2019b) ; multi-level partitioning (Jung and Pramanik, 1996) in G-tree (Zhong et al., 2013, 2015), SPAH (Jung and Pramanik, 1996), CPR (Delling et al., 2017) and (Holzer et al., 2009; Chondrogiannis and Gamper, 2016, 2014). But, their selections are performed in an ad-hoc manner or consider the boundary number, partition balance, and spatial awareness, rather than the SP computation itself.
2.3. Traditional PSP Strategy
The PSP index $L$ typically consists of two components: the partition indexes $\{L_{i}\}$ for each subgraph (partition) $G_{i}$, and the overlay index $\tilde{L}$ for the overlay graph which is composed of the boundary vertices of all partitions, i.e., $L=\{L_{i}\}\cup\tilde{L}$. Regardless of the specific SP algorithm, almost all the existing PSP indexes (Jung and Pramanik, 2002; Delling et al., 2017; Holzer et al., 2009; Chondrogiannis and Gamper, 2014, 2016; Zhong et al., 2013; Li et al., 2019a; Zhong et al., 2015; Ta et al., 2017; Wang et al., 2019, 2016; Liu et al., 2021, 2022b, 2022a; Li et al., 2019c, 2020b; Li et al., 2020a; Zheng et al., 2022; Wang et al., 2021; Farhan et al., 2022; Gubichev et al., 2010; Qiao et al., 2012) construct the PSP index and process the queries under the following procedure.

### v3

@@ Abstract.
Shortest Path (SP) computation is a fundamental operation in many real-life applications such as navigation on road networks, link analysis on social networks, etc. These networks tend to be massive, and graph partitioning is commonly leveraged to scale up the SP algorithms. However, the Partitioned Shortest Path (PSP) index has never been systematically investigated. Moreover, few studies have explored its index maintenance in dynamic networks. In this paper, we survey the dynamic PSP index and propose a universal scheme for its design and analysis. Specifically, we first review the SP algorithms and put forward a novel structure-based partition method classification to facilitate the selection of partition methods. Furthermore, we summarize the existing Pre-boundary PSP strategy and propose two novel strategies (No-boundary and Post-boundary) to improve its index performance. Lastly, we propose a universal scheme with three dimensions (SP algorithm, partition method, and PSP strategy) to facilitate the analysis and design of the PSP index. Based on this scheme, we put forward five new PSP indexes with a prominent query or update efficiency performance. Extensive experiments are conducted to evaluate the performance of the PSP index and the effectiveness of the proposed techniques, with valuable guidance on the PSP index design.
Shortest Path Index, Dynamic Graph, Graph Partition
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Shortest Path (SP) query in networks is an essential building block for various daily life applications. Given an origin-destination pair such as $(s,t)$, the point-to-point shortest path query aims to find the path with the minimum length from vertex $s$ to vertex $t$ on network $G$. In real-life applications, this shortest path could be the path with minimum traveling time in a road network (Ding et al., 2008; Li et al., 2017, 2018, 2019c, 2020b; Bast et al., 2016; Zhou et al., 2025), the fastest connection in the web graph or Internet (Boccaletti et al., 2006; Scheideler, 2002), or the most intimate relationships in the social network (Gong et al., 2016; Kolaczyk et al., 2009; Opsahl et al., 2010). With the development of urban traffic systems and the evolvement of online interaction platforms, real-life networks tend to be massive, which brings great challenges to the scalability of the existing shortest-path algorithms with either heavy memory or expensive search overheads. For example, index-free algorithms such as Dijkstra’s (Dijkstra, 1959) and A* (Hart et al., 1968) directly search on the graphs in a best-first manner for query processing, thus are time-consuming on large networks due to large search space. To this end, index-based algorithms such as Contraction Hierarchy (CH) (Geisberger et al., 2008) and 2-Hop Labeling (HL) (Cohen et al., 2003) are proposed to expedite the query efficiency. In particular, 2-hop labeling (Cohen et al., 2003) such as Pruned Landmark Labeling (PLL) (Akiba et al., 2013) and Hierarchical 2-Hop labeling (H2H) (Ouyang et al., 2018) has the state-of-the-art query efficiency but requires $O(nm^{1/2})$ space to store the index for a graph with $n$ vertices and $m$ edges. Moreover, since the networks are dynamic in nature with evolving edge weights or structures, extra efforts such as space and maintenance overheads are required to handle the updates for shortest path indexes, further deteriorating their scalability.
@@
Graph partitioning is often used to improve the scalability of shortest path algorithms and enable more complicated path problems like time-dependent (Li et al., 2019c, 2020b), constraint (Wang et al., 2016; Liu et al., 2021, 2022), and top-$k$ path enumeration (Yu et al., 2020, 2024). We refer to the shortest path algorithms (indexes) adopting graph partitioning as the Partitioned Shortest Path (PSP) algorithms (indexes) [^1] The term “index” also refers to “index-based algorithm”. We use the terms “algorithm” and “index” interchangeably when the context is clear.. They generally decompose a large network into several smaller ones such that the query processing could be enhanced for index-free algorithms, or the construction time and index size can be reduced for index-based algorithms. As shown in Figure 1, we classify the existing SP algorithms into six categories and compare them briefly in terms of index construction time, storage space, query time, and update time. On one extreme are the direct search algorithms (Dijkstra, 1959; Hart et al., 1968; Li et al., 2020c; Zhang et al., 2020) that require no index and can work in dynamic environments directly but are slow for query answering. Then the partitioned search adds various information to guide and reduce search space (Chondrogiannis and Gamper, 2016; Möhring et al., 2007; Bauer and Delling, 2010) to improve query processing. On the other extreme, index-based solutions such as CH (Geisberger et al., 2008) and HL (Cohen et al., 2003; Akiba et al., 2013; Li et al., 2019b; Ouyang et al., 2018) have larger index sizes, longer construction and maintenance time but faster query performance. Their partitioned versions (Li et al., 2020a, 2019c; Wang et al., 2021; Farhan et al., 2022; Liu et al., 2021, 2023, 2022) are faster to construct with smaller sizes but longer query times. As these works have better performance than their unpartitioned versions, they may create the following misconceptions regarding the benefits of partitions on pathfinding-related problems: 1) Graph partitioning and shortest path algorithms are irrelevant since they are two different research branches. This is often phrased as “graph partitioning is a well-studied problem and their methods could use any of the existing partition solutions” such that only one partition method is used without justification experimentally (Wang et al., 2016; Li et al., 2019c; Liu et al., 2021; Abraham et al., 2011; Chondrogiannis and Gamper, 2014; Yu et al., 2020); 2) Applying graph partitioning is always better, which is the reason why these PSP indexes were proposed in the first place; 3) Different PSP indexes are irrelevant and very different from each other, as evidenced by the fact that they only compared with their non-partitioned counterparts (Chondrogiannis and Gamper, 2016; Li et al., 2020a; Zheng et al., 2022; Liu et al., 2022; Dan et al., 2022) but not with each other. However, these misconceptions are not always true.
Motivation. Although the PSP indexes have been widely applied in the past few decades, they have not been studied systematically. Specifically, the existing solutions typically glue one partition method with one SP algorithm without discriminating their characteristics, figure out a way to query correctly, and then claim superiority. For instance, THop (Li et al., 2019c) utilizes PUNCH (Delling et al., 2011b) for partitions and PLL (Akiba et al., 2013) for boundary and inner indexes, and processed the queries by concatenating inner and boundary indexes. While G-tree (Zhong et al., 2013) uses METIS (Karypis and Kumar, 1998) for hierarchical partitions and builds the distance matrices for vertices on different levels, answering the queries by dynamic programming. These two PSPs seem very different, with different partition methods and path indexes. However, they essentially share the same strategy for organizing index construction and query processing procedures. Besides, most PSP indexes only compare with the non-partitioned counterparts and claim their superiority without comparing with other PSP indexes. Consequently, there lacks a generalized scheme to organize, analyze, and compare the PSP indexes insightfully and fairly. Moreover, how to extend the PSP indexes to dynamic networks is also unknown. To this end, we aim to theoretically and systematically study the dynamic PSP index by conducting a comprehensive survey and proposing a universal scheme to guide its design for different scenarios.
@@
Challenges. However, achieving this goal presents several significant challenges.
Firstly, graph partition methods are normally classified from different perspectives (like vertex-cut / edge-cut, in-memory / streaming, and etc..) with various optimization objectives (such as cut size minimization, balancing, and etc..), but these criteria are not path-oriented and thus hardly help to tell which partition method is suitable for certain path applications. Therefore, we propose a new partitioning classification from the perspective of partition structure for easier partition method selection. We also briefly review the shortest path algorithms used in the PSP index from the perspective of query and update efficiency to guide their selection.
Secondly, it is unclear how to construct, query, and update a new PSP index as all existing works only claim their unique structures can answer queries correctly and seldom involve index maintenance. Therefore, we dive into the PSP strategy, which determines how to build the index, coordinate multiple partitions for correct query answering, and deal with index updates. In particular, we summarize classic pre-boundary PSP strategy and then propose two novel PSP strategies (no-boundary and post-boundary strategies) and index update methods to improve its index performance. Besides, a pruning-based overlay graph optimization is also proposed to improve the index and update efficiency further.
Thirdly, the current PSP indexes are delicately designed structures with tightly coupled components and all claim certain superiority. However, since no effort has been put into the abstract PSP index itself, it is unclear what determines their performance, when to use particular indexes, and how to design a suitable new PSP index for a new scenario. Therefore, based on the rich literature, we propose an insightful and universal PSP index scheme with three critical dimensions, i.e., shortest path algorithm, partition method, and PSP strategy. Based on this scheme, we propose five new PSP indexes to achieve better efficiency in either query or update operations by selecting the specific strategy or method on each dimension as per system requirements.
Contributions. Our contributions and survey structure are listed as follows:
@@
    [Literature Review] We propose a novel structure-based partitioning classification to facilitate the partition method selection for the PSP index, and provide comprehensive literature reviews of corresponding partitioning methods and PSP indexes. Besides, we also briefly review the shortest path algorithms used in PSP index. (Sections 3-4)
@@
    [Improvements] We summarize the traditional PSP strategy as Pre-boundary strategy and propose two novel PSP strategies called No-Boundary and Post-Boundary strategies to improve its index construction and update efficiency. Furthermore, we investigate the index maintenance of these PSP strategies and design a novel pruning-based overlay graph optimization to prune the unnecessary computation. (Section 5)
@@
    [PSP Index Scheme and New PSP Indexes] We propose a universal PSP index scheme by decoupling the PSP index into three dimensions: shortest path algorithm, partition method, and PSP strategy, such that the PSP index can be analyzed and designed systematically. Based on the scheme, we identify and design five new PSP indexes suitable for specific network structures and scenarios to validate the effectiveness of our scheme. (Section 6)
@@
    [Experimental Evaluation] We implement comprehensive experimental studies to explore the performance of various PSP indexes under our proposed universal scheme and provide guidelines for designing PSP indexes with different application scenarios. (Section 7)
2. Preliminary
@@
In this paper, we focus on the dynamic weighted network $G(V,E)$ with $V$ denoting the vertex set and $E$ denoting the edge set. We denote the number of vertices and edges as $n=|V|$ and $m=|E|$. Each edge $e\in E$ is assigned a non-negative weight $w=|e|$, which can increase or decrease in ad-hoc. Note that we focus on edge weight change updates since the structural update (i.e., edge/vertex deletion/insertion) can be converted to edge weight updates and leverage their update mechanisms for structural updates (Ouyang et al., 2020; Zhang et al., 2021c, b, a), as will be discussed in Section 5.3. For each $v\in V$, we represent its neighbors as $N(v)=\{u,|(u,v)\in E\}$. We associate each vertex $v\in V$ with order $r(v)$ indicating its importance in $G$. In this paper, we use boundary-first ordering (Liu et al., 2022) to generate the order for the PSP indexes. For non-partitioned SP indexes, we leverage degree to decide the order of small-world networks while the minimum degree elimination-based ordering (Berry et al., 2003) for the road networks. A path $p=\left<v_{0},v_{1},\dots,v_{k}\right>((v_{i},v_{i+1})\in E,0\leq i<k)$ is a sequence of adjacent vertices with length of $l(p)=\sum_{i=0}^{k-1}w(v_{i},v_{i+1})$, where $w(v_{i},v_{i+1})$ is the edge length from $v_{i}$ to $v_{i+1}$. Given an origin $s$ and destination $t$ vertex (OD pair), the shortest path query $Q(s,t)$ calculates the path with the minimum length between them, i.e., $p_{G}(s,t)$ with the shortest distance $d_{G}(s,t)$. We omit the notation $G$ when the context is clear. Note that we consider $G$ as an undirected graph, and our techniques can be easily extended to directed graphs.
Graph partitioning generally divides a graph $G$ into multiple disjoint subgraphs $\{G_{i}\}$ $(i=1,2,\dots,k)$ with [^2] In this paper, we generally refer to graph partitioning as edge-cut partitioning for easier illustration. $V_{i}\cap V_{j}=\emptyset$ and $\bigcup V_{i}=V$. $\forall v\in G_{i}$, we say $v$ is a boundary vertex if there exists a neighbor of $v$ in the another subgraph, that is $\exists u\in N(v),u\in G_{j}(i\neq j)$. Otherwise, $v$ is an inner vertex. We represent the boundary vertex set of $G_{i}$ as $B_{i}$ and that of $G$ as $B=\bigcup B_{i}$. The PSP indexes then build an overlay graph $\tilde{G}$ on all boundary vertices to connect different partitions and preserve the global distances for all cross-partition queries (i.e., $s\in G_{i},t\notin G_{j},i\neq j$). For $(u,v)\in E$, we say $(u,v)$ is an inter-edge if both its two endpoints $u$ and $v$ are boundary vertices from different subgraphs (i.e., $u\in B_{i},v\in B_{j},i\neq j$). Otherwise, it is an intra-edge. The corresponding edge sets are denoted as $E_{inter}$ and $E_{intra}$, respectively.
Example 0.
Figure 2 presents an example graph $G$ which is partitioned into four subgraphs $G_{1}$ to $G_{4}$, with the red vertices denoting the boundary vertices (e.g., $B_{1}=\{v_{3},v_{10}\}$) and the white vertices representing the inner vertices (e.g., $v_{1}$ and $v_{2}$). Edges outside partitions like $e(v_{3},v_{5})$ are inter-edges, and $e(v_{3},v_{10})$ is an intra-edge. Figure 2-(c) illustrates an overlay graph $\tilde{G}$ built on boundary vertices to connect all subgraphs while preserving the global distances (e.g., $|e_{\tilde{G}}(v_{8},v_{9})|=2$).